If income tax increase by 19% and net income is decrease by 6% then, rate of income tax is


1
Expert's answer
2021-03-12T00:55:29-0500

Income =6 %decrease

Let Income is 100 and income tax is x%

Income income tax Total income

100 x 100 - x

100 19%

1.19x ( 100 - 1.19x )

So, (100 - x) × 94/100 =100 - 1.19x

9400 - 9x = 10000 - 119x

10000 - 9400 = 119x - 9x

600 = 110x

X = 600/ 110 = 5.55%

So income tax rate = 5.55%